Primary 5 question. Michelle had a sum of money. She spent $30 on a dress and 1/5 of the remainder on a skirt. She still had 2/3 of her money left. How much did she have at first? Anwser is $180.
Answer by mananth(16949)

let her have $ x
she spent 30 on dress
Balance she had = x - 30)
She spent 1/5 of it on skirt =0.2
Skirt 0.2 (x- 30 )
Balance she had was (x-30) -0.2(x- 30)
(x-30) -1/5(x- 30) =2x/3
LCD of equation = 15
multiply by 15
15(x-30)-3(x-30)=10x
15x-450-3x+90=10x
12x-10x=360
2x=360
/2
x=180
m.ananth@hotmail.ca